SpaceRef

SpaceRef


NASA Space Station On-Orbit Status 1 December 2004

Status Report From: NASA HQ/SpaceRef/NASA Watch
Posted: Wednesday, December 1, 2004

image SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by SpaceRef.com (copyright © 2004) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below.

After wakeup at the usual time (1:00am EST), the crew s station inspection this morning included the regular monthly routine checkup by FE Sharipov on the DC1 "Pirs" docking module s AZS circuit breakers on the BVP Amp Switch Panel (they should all be On) and the LEDs (light-emitting diodes) of 14 fuses in Fuse Panels BPP-30 & BPP-36.

A major part of today s workload was the reintegration of the Progress 15P cargo ship (TKG) into the Russian segment (RS).  After concluding the obligatory leak check of the SM-to-TKG tunnel/vestibule, the crew opened the transfer hatch and installed the Q/D (quick disconnect) screw clamps (SSVP).  15P was then deactivated and the ventilation air duct installed in the interface passage.  For the purpose of leak check and SSVP install, RS thrusters were temporarily disabled (~9:30am), with the U.S. CMGs remaining in control.   [On 11/24, Progress 15P had been prepared for the eventuality that it may have had to be undocked by TsUP/Moscow via remote control in the unlikely event of an aborted Soyuz redocking last Monday).]

Later in the day, Salizhan Sharipov reinstalled the LKT local temperature sensor commutator (TA251M1B) of the BITS2-12 onboard telemetry system and its ROM (read-only memory) unit (TA765B) in the cargo ship.  The LKT was subsequently switched on by the ground to complete the basic configuration.   [The US-21 matching unit with its associated commutator gear (which provide the electronic interface between the SM and the TKG for SM computer control of the Progress thrusters) will be reinstalled tomorrow.]

CDR Chiao conducted two inventory audits to update the books on currently available supplies and their actual condition.  The first review focused on computer hardware/spares, such as hard disk drives (HDDs), by auditing and consolidating the contents of the various onboard computer bags.  There are about three hours scheduled this week, with any necessary runover to be added to the job jar task list.

The second audit dealt with the contents of the AMP (ambulatory medical pack), to ensure that all available medications are in their proper locations and no pills are crushed.   [The two-hour task was supported by an uplinked reference list of AMP drugs, with both generic names and brand names, correlated for unambiguous identification.]

Leroy filled out his regular weekly FFQ (food frequency questionnaire), his fifth, that keeps a log of his nutritional intake over time on special MEC (medical equipment computer) software.   [The FFQ records amounts consumed during the past week of such food items as beverages, cereals, grains, eggs, breads, snacks, sweets, fruit, beans, soup, vegetables, dairy, fish, meat, chicken, sauces & spreads, and vitamins.]

Work continues on the ground to put together the food manifest for the 12/24 launch of Progress 16P.  A third audit by the CDR, tallying the food provisions in the pantry (Russian: SOP), is scheduled for tomorrow to actualize the IMS (inventory management system) on the shrinking supply of food rations.  

In the SM, Salizhan started the regeneration cycle on absorbent bed #2 of the BMP harmful impurities unit and switched channel #1 from its 24-hr regeneration process back to Purify mode.  Because of this operation, the Elektron remains powered down until tomorrow (~5:10am).   [The "bakeout" cycle for the filter beds is repeated every 20 days.  Each bakeout to space vacuum takes about 24 hours. Re Elektron: In order to evaluate the feasibility of using a specific plug for capping the AVK-5 hydrogen vacuum vent nozzle during the Electron-VM machine s H2 vent valve (ZLVK) replacement, TsUP/Moscow has requested U.S. support in considering external photography of the vent nozzle area with maximum possible resolution.  The nozzle is located on the conical part of the SM Working Compartment (RO) and would apparently have to be capped via EVA for the valve R&R.]

Chiao conducted today s daily routine maintenance of the SOZh environment control & life support systems in the SM, including ASU toilet subsystems.

Previous Reports

ISS On-orbit Status [HQ]
ISS Status [JSC]
Shuttle Processing [KSC]

The crew performed their daily 2.5-hr. physical exercise program on TVIS treadmill, RED exerciser, CEVIS cycle and VELO cycle with bungee cord load trainer.  During his one-hour workout on the CEVIS, Leroy was to wear his heart rate watch along with the chest strap.

Afterwards, the CDR transferred the daily TVIS and RED exercise data files to the MEC for downlink, as well as the daily wristband HRM (heart rate monitor) data, followed by their erasure on the HRM storage medium.

Overnight, TsUP started another extensive test run on the SM s ASN-2401 antenna hardware of the Russian ASN-M satellite navigation system, this time involving the combined system.   [When finally operational, ASN will use GLONASS satellites (the Russian GPS equivalent) to update the state vector (SV, position & velocity data) without using the ground (which up to now has to uplink daily SV updates) or requiring SV transfers from the U.S. segment (USOS) from time to time.]

On Monday (11/29), antenna assembly #4 (AA-4) of the ISS GPS equipment dropped its usable power (expressed as signal-to-noise [S/N] ratio) to a level rendering it unusable for measurements for the U.S. SIGIs (Space Integrated GPS/Inertial Navigation Systems).  A similar signature was observed on the failed #2 antenna (e.g., see 11/17 Status).  With only two of four GPS antennas operational, SIGI can no longer provide station attitude data.  State vector determinations, however, require only one AA.   [The determination of whether either antenna is failed is not yet final.  As investigation continues, the U.S. attitude system has been reset to accept Russian data as needed.]

The crew has located the solid shim for the EMU (extravehicular mobility unit) 3005 water pump that went missing during the 11/10 attempt to replace the pump rotor.  Planning is underway to resume the EMU cooling system s repair at the earliest opportunity, probably in two weeks.   [Would you believe how it was found?  Sharipov happened to see the tiny metal ring float by as he was in the DC1 to set up for a teleconference with TsUP.  The shim was reported to be in good condition, and pictures were sent down to the ground for evaluation.]

Today's CEO photo targets, no longer limited in the current LVLH attitude, were Salamat Basin fans, Chad (weather was clear for high-resolution mapping swaths across this megafan complex.  Looking to the right of track for divergent drainage patterns and wetlands that mark fan boundaries), Internal Waves. Caribbean Sea (this overpass provided an opportunity for internal wave photography.  The sunglint point was located to the right of track), and Palmerston Island reef, central S Pacific (weather was predicted to be clear for high-resolution nadir photography of the reef.  Mapping swaths across the reef are useful for assessment of bleaching and changes to reef morphology.

CEO photography can be viewed and studied at the websites:

See also the website "Space Station Challenge" at:

To view the latest photos taken by the expedition 10 crew visit:

Expedition 10 Flight Crew Plans can be found at http://spaceflight.nasa.gov/station/timelines/

Upcoming Key Events: 

  • Progress 15P undocking & destructive reentry -- 12/23;
  • Progress 16P launch -- 12/24;
  • Progress 16B docking -- 12/26;
  • EVA-12 -- January 2005 (1/27?).

ISS Location NOW

Full Size/Update
Real Time ISS Tracker - More Links

ISS Orbit  (as of this morning, 5:44am EST [= epoch]):

  • Mean altitude -- 356.3 km
  • Apogee height -- 359.2 km
  • Perigee height -- 353.4 km
  • Period -- 91.67 min.
  • Inclination (to Equator) -- 51.64 deg
  • Eccentricity -- 0.000428
  • Solar Beta Angle -- -56.5 deg (magnitude increasing)
  • Orbits per 24-hr. day -- 15.71
  • Mean altitude loss in last 24 hours -- 50 m
  • Revolutions since FGB/Zarya launch (Nov. 98) -- 34459

 

ISS Altitude History

Apogee height -- Mean Altitude -- Perigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html. In addition, information on International Space Station sighting opportunities can be found at http://spaceflight.nasa.gov/realdata/sightings/ on NASA's Human Spaceflight website. The current location of the International Space Station can be found at http://science.nasa.gov/temp/StationLoc.html at NASA's Marshall Space Flight Center. Additional satellite tracking resources can be found at http://www.spaceref.com/iss/tracking.html.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

NanoRacks